OpenStudy (anonymous):

3x+1/2+x-1/3=15/2 X=?

OpenStudy (anonymous):

here it is

OpenStudy (anonymous):

thanks broo berry good

hero (hero):

Better method: \(\frac{3x + 1}{2} + \frac{x -1}{3} = \frac{15}{2} \) \( \frac{x -1}{3} = \frac{15}{2} - \frac{3x + 1}{2} \) \( \frac{x -1}{3} = \frac{15 - 3x - 1}{2}\) \( \frac{x -1}{3} = \frac{14 - 3x}{2}\) \(\small 2(x-1) = 3(14 - 3x)\) \(\small 2x - 2 = 42 - 9x\) \(\small 9x + 2x = 42 + 2\) \(\small 11x = 44\) \(\small x = \frac{44}{11}\) \(\small x = 4\)